# Hypergraph Foundation Model

### v1

@@ Abstract
Hypergraph neural networks (HGNNs) effectively model complex high-order relationships in domains like protein interactions and social networks by connecting multiple vertices through hyperedges, enhancing modeling capabilities, and reducing information loss. Developing foundation models for hypergraphs is challenging due to their distinct data, which includes both vertex features and intricate structural information. We present Hyper-FM, a Hypergraph Foundation Model for multi-domain knowledge extraction, featuring Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding for vertex feature representation and Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction for structural information. Additionally, we curate 10 text-attributed hypergraph datasets to advance research between HGNNs and LLMs. Experiments on these datasets show that Hyper-FM outperforms baseline methods by approximately 13.3%, validating our approach. Furthermore, we propose the first scaling law for hypergraph foundation models, demonstrating that increasing domain diversity significantly enhances performance, unlike merely augmenting vertex and hyperedge counts. This underscores the critical role of domain diversity in scaling hypergraph models.
Index Terms:
Hypergraph Neural Networks, Foundation Model, High-Order Learning, Hypergraph Learning.
@@ I Introduction
Regarding the extraction of knowledge from hypergraph vertex features, existing hypergraph models predominantly utilize Bag-of-Words (BoW) [16] representations. This approach leads to ever-increasing feature dimensions as textual descriptions grow, and varying vertex feature dimensions across different domains complicate the alignment of vertex features. Current methods attempt to extract fixed-dimensional embeddings from raw textual data using techniques like Singular Value Decomposition (SVD) [17] or Language Models (LM) [12, 13, 14, 15]. However, these approaches often overlook structural information, resulting in suboptimal feature representations. For instance, vertices with identical feature representations but located in different positions within the hypergraph should possess distinct semantic embeddings. Directly using LMs for feature extraction fails to incorporate structural information, raising the question: how can structural information be injected into LM-extracted features to enhance their discriminative power?
In terms of extracting structural knowledge from hypergraphs, the most straightforward approach is to pre-train [18, 19, 20] independently on different domain-specific hypergraph structures to generate a hypergraph foundation model. However, this approach yields poor results, as shown in Fig. 1. Pre-training on source domain hypergraph structures and then applying the model to target domain downstream tasks leads to negative transfer. This occurs because relational data from different domains exhibit significant structural pattern differences. Consequently, how can multi-domain knowledge extraction and transfer be effectively achieved? Existing methods typically concatenate structures directly, but we observe that this approach only performs adequately on small-scale datasets. It remains unclear whether large-scale knowledge extraction and transfer are feasible. Additionally, direct concatenation can result in the blending of domain-specific knowledge, as evidenced by the performance bottleneck observed in the experiments presented in Tab. VI. Thus, how can knowledge be transferred across domains while preserving the unique semantic information of each domain?
@@
Furthermore, we have manually curated and constructed 10 TAHG datasets to facilitate cross-disciplinary research between HGNNs and LLMs. Experiments on these 10 hypergraph datasets demonstrate that our proposed foundation model outperforms baseline methods by an average of approximately 13.3%, validating its effectiveness. Additionally, we are the first to propose scaling laws for hypergraph computation. Our findings reveal that, unlike other domains, simply increasing the scale of data in terms of vertex and hyperedge counts does not enhance the foundational model’s power, as illustrated in Fig. 5. However, increasing the number of domains—thereby exposing the model to a wider variety of relational structures—significantly enhances the model’s performance. This suggests a promising direction for future development. In summary, our contributions are threefold:
@@
    We curate and release 10 text-attributed hypergraph (TAHG) datasets, fostering the integration of hypergraph computation and language model research.
@@
    We introduce the first hypergraph foundation model for extracting vertex and structural knowledge from multi-domain hypergraph data. Leveraging hierarchical representations and a sampling-based Hierarchical Multi-Hypergraph, our model achieves an average performance improvement of approximately 13.3% across 10 TAHG datasets, demonstrating its effectiveness.
@@ II Related works
II-A Foundataion Models
Foundation models have significantly transformed both NLP and CV domains. In NLP, BERT [12], introduced by Devlin et al. revolutionized context-aware embeddings by utilizing a bidirectional training approach, enhancing performance on tasks like question answering. Following this, GPT-2 [14] and GPT-3 [15] from OpenAI showcased the power of generative pre-training, with GPT-3’s 175 billion parameters demonstrating impressive few-shot learning capabilities across various applications. T5 [24] further unified NLP tasks under a text-to-text framework, simplifying the training process and achieving strong results in diverse tasks. RoBERTa [13], a robustly optimized version of BERT, improved training efficiency by removing the next-sentence prediction objective and leveraging larger datasets. In CV, Models like DINO [25] utilized self-supervised learning methods, leveraging knowledge distillation to learn rich visual representations without relying on labeled data. SAM [26] enhanced image segmentation capabilities, allowing for precise object delineation, while SegGPT [27] merged segmentation and generative models to tackle complex visual tasks. Lastly, Visual ChatGPT [28] combined the strengths of GPT-3 [15] and visual understanding, enabling interactive applications that integrate text and image processing. Together, these developments highlight the significant impact of foundation models in advancing machine learning methodologies across diverse applications.
@@ II-B Graph Foundation Models
In recent years, several innovative graph foundation models have emerged, significantly advancing the field of graph data processing. InstructGLM [29] is a generative language model specifically designed for graph structures, enabling users to create and modify graphs through natural language instructions, thereby enhancing interactivity with graph data. GraphGPT [30] combines graph neural networks with generative pre-trained transformers, focusing on generating and reasoning about graph data, which facilitates tasks such as graph description and generation through contextual information. LLaGA [31] aligns language models with graph structures to improve representation learning by jointly training on both data types, boosting performance on tasks like node classification and link prediction. Finally, GALLM [32] integrates graph information into language models, strengthening the connection between generated text and graph data, making it suitable for applications like knowledge graph construction and information retrieval.
@@ II-C Hypergraph Neural Networks
The development of Hypergraph Neural Networks (HGNNs) has significantly advanced the modeling of complex relationships in graph-structured data. At the beginning, HGNN [33] introduced hypergraph convolution operations that effectively capture higher-order relationships among nodes, enhancing the representation capabilities for graph-structured data. Based on this, HGNN+ [34] incorporated a more sophisticated neighbor aggregation mechanism, using multilevel information fusion to improve node feature learning. HyperGCN [35] further advanced this field by implementing hypergraph convolution layers for efficient information propagation directly on hypergraphs, demonstrating robust performance across various tasks. The HNHN [36] model adopted a hierarchical node representation approach, enabling layer-wise information dissemination through multilayer hypergraph structures, thus enhancing modeling capacity for intricate relationships. In contrast, HGAT [37] combined self-attention mechanisms with hypergraph structures, allowing dynamic adjustment of the aggregation weights to automatically learn the importance of different nodes and hyperedges. UniGNN [38] proposed a unified framework that integrates multiple graph structures, facilitating the handling of diverse graph data types and enhancing model flexibility. Furthermore, HyperSAGE [39] draws inspiration from GraphSAGE [40], devising a sampling-based aggregation strategy to reduce computational complexity and improve scalability for large hypergraphs. Lastly, AllSet [41] introduced concepts from set theory to optimize hypergraph representation, thereby increasing the model’s capacity for understanding and reasoning about complex relationships.
@@ III Preliminaries
A hypergraph can be represented as ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}=\{v_{1},v_{2},\ldots\}$ is the set of vertices and =${\mathcal{E}}=\{e_{1},e_{2},\ldots\}$ is the set of hyperedges. Typically, each vertex is associated with some textual descriptions, denoted as ${\mathcal{T}}=\{T_{1},T_{2},\ldots\}$, where $T_{i}$ is the textual description for vertex $v_{i}$.
@@
Hypergraph Foundation Models.
@@
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{({\mathcal{T}}_{i},{\mathcal{G}}_{i})\in{\mathcal{D}}^{s}}{\mathcal{L}}_{\text{pre}}\left(f({\mathcal{T}}_{i},{\mathcal{G}}_{i}\mid\theta)\right),$ (2)
where $\theta^{*}$ denotes the optimized parameters of the Hypergraph Foundation Model after pre-training. The dataset ${\mathcal{D}}^{s}=\{({\mathcal{T}}_{1},{\mathcal{G}}_{1}),({\mathcal{T}}_{2},{\mathcal{G}}_{2}),\ldots\}$ comprises multi-domain hypergraph data, with each pair $({\mathcal{T}}_{i},{\mathcal{G}}_{i})$ containing the textual attributes ${\mathcal{T}}_{i}$ of the vertices and the structural information ${\mathcal{G}}_{i}$ of the hypergraph. The textual attributes ${\mathcal{T}}_{i}$ provide rich semantic information that complements the structural data captured by ${\mathcal{G}}_{i}$. The loss function ${\mathcal{L}}_{\text{pre}}$ corresponds to a predefined self-supervised hypergraph task designed to encode knowledge from multi-domain hypergraphs into the parameters. The function $f_{\theta}$ represents the hypergraph neural network model with trainable parameter $\theta$. Through self-supervised tasks, the model captures and integrates knowledge from diverse hypergraph domains into the foundational parameters $\theta^{*}$.
@@
$\theta^{\prime}=&\operatorname*{arg\,min}_{\theta^{\prime}}\sum_{(T_{i},Y_{i})\in({\mathcal{T}}^{t},{\mathcal{Y}}^{t})}{\mathcal{L}}_{\text{down}}\left(p_{\theta}(T_{i},{\mathcal{G}}^{t}),Y^{t}\right)\\ &\text{subject to}\quad\theta^{\prime(0)}\leftarrow\theta^{*}\end{aligned},$ (3)
where $({\mathcal{T}}^{t},{\mathcal{G}}^{t},{\mathcal{Y}}^{t})$ represents the target domain hypergraph data, including textual vertex attributes ${\mathcal{T}}^{t}$, hypergraph structure ${\mathcal{G}}^{t}$, and vertex labels ${\mathcal{Y}}^{t}$. The textual attributes ${\mathcal{T}}^{t}$ provide semantic context for the vertices, while the structural information ${\mathcal{G}}^{t}$ captures the relationships between vertices via hyperedges. The label set ${\mathcal{Y}}^{t}=\{Y_{1},Y_{2},\ldots,Y_{N}\}$ consists of labels for each vertex in the target hypergraph, essential for supervised fine-tuning. The loss function ${\mathcal{L}}_{\text{down}}$ pertains to the downstream vertex classification task, aiming to minimize the discrepancy between the predicted labels $p_{\theta}({\mathcal{T}}_{i},{\mathcal{G}}^{t})$ and the ground true labels $Y^{t}$. The parameter $\theta^{\prime}$ represents the fine-tuned parameters of the hypergraph neural network tailored to the target domain. The predictive function $p_{\theta}({\mathcal{T}}_{i},{\mathcal{G}}^{t})$ maps the input textual attributes and hypergraph structures to label probabilities using the model parameters $\theta$. The initialization condition $\theta^{\prime(0)}\leftarrow\theta^{*}$ ensures that fine-tuning begins with the pre-trained foundation parameters, facilitating the transfer of multi-domain knowledge to the specific task.
@@ IV Methodology
In this section, we present the framework for Multi-Domain Hypergraph Foundation Models (Hyper-FM). We begin with an overview of the proposed framework, highlighting its design and primary objectives. The framework comprises two core modules: Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding, which generates structure-aware vertex embeddings by capturing high-order neighborhood information, and Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction, which extracts and integrates structural knowledge from multiple hypergraph domains using a hierarchical sampling strategy. We then describe the pre-training process on multi-domain hypergraph datasets through self-supervised learning and outline the strategies for adapting the pre-trained foundation model to specific domain hypergraphs via fine-tuning. Additionally, we also provide comparisons with several classic graph neural network foundation model frameworks to illustrate the advantages of our approach.
@@ IV-B1 High-Order Neighborhood Vectorization
where ${\mathcal{D}}_{\text{nbr}}(v)_{[i]}=1$ if vertex $v$ is connected to vertex $i$ via a hyperedge, and 00 otherwise. Importantly, if two vertices $v_{i}$ and $v_{j}$ share identical neighborhoods, their vectors satisfy ${\mathcal{D}}_{\text{nbr}}(v_{i})={\mathcal{D}}_{\text{nbr}}(v_{j})$. Leveraging this vectorization, we define the training objective for the LM as follows:
@@
IV-B2 Hierachical Neighborhood Representation

### v2

@@ Abstract
Hypergraph neural networks (HGNNs) effectively model complex high-order relationships in domains like protein interactions and social networks by connecting multiple vertices through hyperedges, enhancing modeling capabilities, and reducing information loss. Developing foundation models for hypergraphs is challenging due to their distinct data, which includes both vertex features and intricate structural information. We present Hyper-FM, a Hypergraph Foundation Model for multi-domain knowledge extraction, featuring Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding for vertex feature representation and Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction for structural information. Additionally, we curate 11 text-attributed hypergraph datasets to advance research between HGNNs and LLMs. Experiments on these datasets show that Hyper-FM outperforms baseline methods by approximately 13.4%, validating our approach. Furthermore, we propose the first scaling law for hypergraph foundation models, demonstrating that increasing domain diversity significantly enhances performance, unlike merely augmenting vertex and hyperedge counts. This underscores the critical role of domain diversity in scaling hypergraph models.
@@ I Introduction
Regarding the extraction of knowledge from vertex features, existing hypergraph models predominantly utilize Bag-of-Words (BoW) [16] representations. This approach leads to ever-increasing feature dimensions as textual descriptions grow, and varying vertex feature dimensions across different domains complicate the alignment of vertex features. Current methods attempt to extract fixed-dimensional embeddings from raw textual data using techniques like Singular Value Decomposition (SVD) [17] or Language Models (LM) [12, 13, 14, 15]. However, these approaches often overlook structural information, resulting in suboptimal feature representations. For instance, vertices with identical feature representations but located in different positions within the hypergraph should possess distinct semantic embeddings. Directly using LMs for feature extraction fails to incorporate structural information, raising the question: how can structural information be injected into LM-extracted features to enhance their discriminative power?
In terms of extracting structural knowledge from hypergraphs, the most straightforward approach is to pre-train [18, 19, 20] independently on different domain-specific hypergraph structures to generate a hypergraph foundation model.
However, this approach yields poor results, as shown in Fig. 1. Pre-training on source domain hypergraph structures and then applying the model to target domain downstream tasks leads to negative transfer. This occurs because relational data from different domains exhibit significant structural pattern differences. Specifically, the lower part of Fig. 1 illustrates the underlying reason for performance degradation when transferring from Aminer-Text to other datasets. We plot the statistical distributions of hyperedge degrees and observe that Movielens-Text shares a distribution pattern similar to Aminer-Text, resulting in only a marginal performance drop of 0.64%. In contrast, as the distributional discrepancy increases, the performance degradation becomes more pronounced. For instance, the hypergraph distribution of PPI-Text differs substantially from that of Aminer-Text, with distinct peak positions and fluctuation patterns, leading to a significant performance drop of 18.07%. Consequently, how can multi-domain knowledge extraction and transfer be effectively achieved?
Existing methods typically concatenate structures directly, but we observe that this approach only performs adequately on small-scale datasets. It remains unclear whether large-scale knowledge extraction and transfer are feasible. Additionally, direct concatenation can result in the blending of domain-specific knowledge, as evidenced by the performance bottleneck observed in the experiments presented in Tab. VI. Thus, how can knowledge be transferred across domains while preserving the unique semantic information of each domain?
@@
Furthermore, we have manually curated and constructed 11 TAHG datasets to facilitate cross-disciplinary research between HGNNs and LLMs. Experiments on these 11 hypergraph datasets demonstrate that our proposed foundation model outperforms baseline methods by an average of approximately 13.4%, validating its effectiveness. Additionally, we are the first to propose scaling laws for hypergraph computation. Our findings reveal that, unlike other domains, simply increasing the scale of data in terms of vertex and hyperedge counts does not enhance the foundational model’s power, as illustrated in Fig. 5. However, increasing the number of domains—thereby exposing the model to a wider variety of relational structures—significantly enhances the model’s performance. This suggests a promising direction for future development. In summary, our contributions are threefold:
@@
    We curate and release 11 text-attributed hypergraph (TAHG) datasets, fostering the integration of hypergraph computation and language model research.
@@
    We introduce the first hypergraph foundation model for extracting vertex and structural knowledge from multi-domain hypergraph data. Leveraging hierarchical representations and a sampling-based Hierarchical Multi-Hypergraph, our model achieves an average performance improvement of approximately 13.4% across 11 TAHG datasets, demonstrating its effectiveness.
@@ II Related works
II-A Foundation Models
Foundation models have profoundly transformed NLP and CV. In NLP, BERT [12] revolutionized context-aware embeddings through bidirectional training, boosting tasks like question answering. GPT-2 [14] and GPT-3 [15] further demonstrated the power of generative pre-training, with GPT-3’s 175B parameters enabling impressive few-shot learning. T5 [24] unified NLP tasks under a text-to-text framework, while RoBERTa [13] improved training efficiency by removing the next-sentence prediction objective and using larger datasets. In CV, models like DINO [25] leveraged self-supervised learning and knowledge distillation to learn rich visual representations without labels. SAM [26] advanced image segmentation, SegGPT [27] integrated segmentation and generation for complex tasks, and Visual ChatGPT [28] combined GPT-3 with visual understanding for interactive multimodal applications. Collectively, these advances underscore the broad impact of foundation models in modern machine learning.
@@ II-B Graph Foundation Models
In recent years, several innovative graph foundation models have emerged, significantly advancing the field of graph data processing. InstructGLM [29] is a generative language model specifically designed for graph structures, enabling users to create and modify graphs through natural language instructions, thereby enhancing interactivity with graph data. GraphGPT [30] combines graph neural networks with generative pre-trained transformers, focusing on generating and reasoning about graph data, which facilitates tasks such as graph description and generation through contextual information. LLaGA [31] aligns language models with graph structures to improve representation learning by jointly training on both data types, boosting performance on tasks like vertex classification and link prediction. Finally, GALLM [32] integrates graph information into language models, strengthening the connection between generated text and graph data, making it suitable for applications like knowledge graph construction and information retrieval.
@@ II-C Hypergraph Neural Networks
Hypergraph Neural Networks (HGNNs) have greatly advanced the modeling of high-order relationships in graph-structured data. HGNN [33] first introduced hypergraph convolution to capture higher-order vertex relations, while HGNN+ [34] enhanced neighbor aggregation via multilevel information fusion. HyperGCN [35] implemented efficient hypergraph convolution layers, and HNHN [36] adopted hierarchical vertex representations for multilayer information propagation. HGAT [37] integrated self-attention with hypergraphs to dynamically learn vertex and hyperedge importance. UniGNN [38] provided a unified framework for diverse graph structures, improving model flexibility. HyperSAGE [39], inspired by GraphSAGE [40], introduced sampling-based aggregation for scalability. AllSet [41] leveraged set theory to optimize hypergraph representation, and DHHNN [42] is a variational autoencoder-based framework for multimodal data integration and vertex classification, which combines hyperbolic geometry, dynamic hypergraphs and self-attention to address the static structure limitation of existing models.
II-D Pioneer in Hypergraph Foundation Models
Recent advances have seen pioneering hypergraph foundation models emerge, expanding hypergraph learning across domains. HGFM [43] is a high-order correlation-driven model for brain disease diagnosis, learning an encoder via self-supervised pretraining on high-order structures and adapting to downstream tasks through few-shot fine-tuning. HYPER [44] is a foundation model for inductive link prediction in knowledge hypergraphs, generalizing to novel entities and relations by encoding entities with their hyperedge positions. DMCL-HFM [45] targets whole slide images, using vertex-hyperedge dual-masked modeling and contrastive learning for pretraining to integrate local and global dependencies. The hypergraph-driven landmark detection model [46] supports cardiac function quantification, leveraging an adaptive dynamic hypergraph backbone and bidirectional hypergraph spatio-temporal decoder. Notably, these models are tailored to specific domain tasks (e.g., brain disease diagnosis), highlighting the lack of a universal hypergraph foundation model framework for multiple scenarios.
@@ III Preliminaries
A hypergraph can be represented as ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}=\{v_{1},v_{2},\ldots\}$ is the set of vertices and ${\mathcal{E}}=\{e_{1},e_{2},\ldots\}$ is the set of hyperedges. Typically, each vertex is associated with some textual descriptions, denoted as ${\mathcal{T}}=\{T_{1},T_{2},\ldots\}$, where $T_{i}$ is the textual description for vertex $v_{i}$.
@@
Hypergraph Foundation Models
@@
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{({\mathcal{T}}_{x},{\mathcal{G}}_{x})\in{\mathcal{D}}^{\text{src}}}{\mathcal{L}}_{\text{pre}}\left(f({\mathcal{T}}_{x},{\mathcal{G}}_{x}\mid\theta)\right),$ (2)
where $\theta^{*}$ denotes the optimized parameters of the Hypergraph Foundation Model after pre-training. The dataset ${\mathcal{D}}^{\text{src}}=\{({\mathcal{T}}_{1},{\mathcal{G}}_{1}),({\mathcal{T}}_{2},{\mathcal{G}}_{2}),\ldots\}$ comprises multi-domain hypergraph data, with each pair $({\mathcal{T}}_{x},{\mathcal{G}}_{x})$ containing the textual attributes ${\mathcal{T}}_{x}$ of the vertices and the structural information ${\mathcal{G}}_{x}$ of the hypergraph. The textual attributes ${\mathcal{T}}_{x}$ provide rich semantic information that complements the structural data captured by ${\mathcal{G}}_{x}$. The loss function ${\mathcal{L}}_{\text{pre}}$ corresponds to a predefined self-supervised hypergraph task designed to encode knowledge from multi-domain hypergraphs into the parameters. The function $f_{\theta}$ represents the hypergraph neural network model with trainable parameter $\theta$. Through self-supervised tasks, the model captures and integrates knowledge from diverse hypergraph domains into the foundational parameters $\theta^{*}$.
@@
$\begin{aligned} \theta^{\prime}=&\operatorname*{arg\,min}_{\theta^{\prime}}\sum_{(T_{x},Y_{x})\in({\mathcal{T}}^{\text{tar}},{\mathcal{Y}}^{\text{tar}})}{\mathcal{L}}_{\text{down}}\left(p_{\theta}(T_{x},{\mathcal{G}}^{\text{tar}}),Y_{x}\right)\\ &\text{subject to}\quad\theta^{\prime(0)}\leftarrow\theta^{*}\end{aligned},$ (3)
where $({\mathcal{T}}^{\text{tar}},{\mathcal{G}}^{\text{tar}},{\mathcal{Y}}^{\text{tar}})$ represents the target domain hypergraph data, including textual vertex attributes ${\mathcal{T}}^{\text{tar}}$, hypergraph structure ${\mathcal{G}}^{\text{tar}}$, and vertex labels ${\mathcal{Y}}^{\text{tar}}$. The textual attributes ${\mathcal{T}}^{\text{tar}}$ provide semantic context for the vertices, while the structural information ${\mathcal{G}}^{\text{tar}}$ captures the relationships between vertices via hyperedges. The label set ${\mathcal{Y}}^{\text{tar}}=\{Y_{1},Y_{2},\ldots,Y_{N}\}$ consists of labels for each vertex in the target hypergraph, essential for supervised fine-tuning. The loss function ${\mathcal{L}}_{\text{down}}$ pertains to the downstream vertex classification task, aiming to minimize the discrepancy between the predicted labels $p_{\theta}(T_{x},{\mathcal{G}}^{\text{tar}})$ and the ground true labels $Y^{t}$. The parameter $\theta^{\prime}$ represents the fine-tuned parameters of the hypergraph neural network tailored to the target domain. The predictive function $p_{\theta}(T_{x},{\mathcal{G}}^{\text{tar}})$ maps the input textual attributes and hypergraph structures to label probabilities using the model parameters $\theta$. The initialization condition $\theta^{\prime(0)}\leftarrow\theta^{*}$ ensures that fine-tuning begins with the pre-trained foundation parameters, facilitating the transfer of multi-domain knowledge to the specific task.
@@ IV Methodology
In this section, we present the Multi-Domain Hypergraph Foundation Model (Hyper-FM). First, we overview its design and core objectives. Hyper-FM has two core modules: Hierarchical High-Order Neighbor Guided Vertex Knowledge Embedding (generates structure-aware vertex embeddings) and Hierarchical Multi-Hypergraph Guided Structural Knowledge Extraction (integrates cross-domain structural knowledge). We then describe its self-supervised multi-domain pre-training and domain-specific fine-tuning, and compare it with classic graph foundation models to show its advantages.
@@ IV-B1 High-Order Neighborhood Vectorization
where ${\mathcal{D}}_{\text{nbr}}(v)_{[i]}=1$ if vertex $v$ is connected to vertex $i$ via a hyperedge, and 0 otherwise. Importantly, if two vertices $v_{i}$ and $v_{j}$ share identical neighborhoods, their vectors satisfy ${\mathcal{D}}_{\text{nbr}}(v_{i})={\mathcal{D}}_{\text{nbr}}(v_{j})$. Leveraging this vectorization, we define the training objective for the LM as follows:
@@
IV-B2 Hierarchical Neighborhood Representation